Systems Staff Meeting Notes 3/1/99

Systems Staff Meeting Notes
March 1, 1998

Present: Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Review Minutes of 2/22/99
  1. RIR#1 - phantom problems in Preservation have not recurred, switch installation seems to have cured the problem.
Ad Hoc
  1. NTAC meeting last Friday, Feb. 26 - discussion on the Network Security Scans. General opinion was that the inspectors weren't really up on the interpretation of the scans. Also, their scheduling of the scans left a lot to be desired. Maintenance windows were another topic - would try to keep them outside of regular business hours. Opposition to the windows basically dictates that no time is a good time. Windows are now defined as: Tuesday 5-7AM, Sunday 5-10AM - if there is a major problem they will back off in time to correct it. The previous windows had an unscheduled or implied hour after the allotted time for backing off if necessary. Notification should occur by noon the day prior to the maintenance.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. We surveyed out some old equipment and received some more from various departments.
  2. 11 Okidata laser printers have been received.
  3. 20 Cows have been received. There has been no distribution list generated yet. 20 more Cows are on order.
  4. Prices are down - $1500 for a 400-450mhz machine; 17 or 19 Gig drives are going for $300.
  5. Bobbie is requesting the installation of another phone for the work area desk across from the students.
Networking
  1. Wiring project- David from Global and Judy at Digital are both awaiting prices for cabinets (some departments want the routers closed-in to reduce the noise level). Both will give an estimated cost in their bids.
  2. Need to update the software on the CISCO switch for Preservation. Will need to go into the 1st floor Com closet, TFTP via laptop - need copy - and it will create a short period of down time.
  3. Smathers NT6 - some folder permissions have been changed for Education. Judiaca was trying to access a temporary directory & was unable.
  4. Smathers NT7 permissions - mostly for SQL server. NT7 files were moved to NT2. Bill is working on retrofitting domain groups on the server.
  5. Bill will try to create an unsecured folder on each server.
  6. Quotas are being reached - 30M mailbox limits - people need to clean off space, but we will inevitably have to expand it. The state requires that we keep all email that we send for 3 years.
  7. Do not use the Web server from a direct internet connection on the same feed for mail - it has a security hole, user in domain can gain access.
Software Projects Status
  1. Mellon is in Erich's hands, complete with changes. Nelda is working on the user documentation.
  2. Winston met with Alena Aissing, she brought the database from her home computer - it probably is not good to run the first InterDev test with it.
  3. Ran invoicing for February - restored from backups.
  4. Installed new NT machine for David Hickey so he can run Japanese, Chinese & Korean programs through OCLC.
  5. Winston has received a lot of virus calls lately - he urges all to keep the virus checker updated. Most of the viruses are the macro type. Erika actually had the Happy 99 virus on her computer and Winston had to remove it.
  6. LASSIE - Rich, Barbara & Marie are connected. Brian decided not to delete dummy records yet. Working on web version - what has to be done for each particular unit. Needs access to web space - individual folders would be good. Suzy notes that folder & file names can have no spaces or squiggles.
  7. Ying is working on the IR database. She has devised 3 functions: software management, hardware management & department information. Need to add more vendor information to the database. Bill found many 'Batc' for model names - these are incorrect. Classes need to be revamped.
  8. ATC - need to get going on this soon, plan on beta testing in June. FCLA will make the CD, we need to specify the format.
DP Coordinator
  1. The third batch of yearly Compustat tapes arrived on Friday - in ASCII format. Bobbie will have to convert the program - this is high priority.
  2. Students will be on Spring Break next week.
  3. Termites - a specific time has not been scheduled yet for the eradication. The tiles are asbestos, workers may show up in moon suits.
Calendar
  1. Debra out Friday, March 4.
  2. St. Patrick's Day - Wednesday, March 17.


Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu